A constant force F acts on a truck over a distance s and for a time t. The momentum gained by the truck is?
1
Expert's answer
2020-11-23T10:29:41-0500

Explanations & Calculations

  • According to the equation "\small F =\large \frac{\Delta mv}t{}" this question could be solved

"\qquad\qquad
\begin{aligned}
\small F &= \small \frac{\Delta mv}{t}\\
\small \Delta mv&= \small F\times t\\
\end{aligned}"

  • Assuming no friction, the change in the momentum (what is gained by an object over the time) is given by "\small F\times t" for a constant force.